OpenStudy (anonymous):

How do I simplify (1 + tan x)^2? Thanks.

Directrix (directrix):

You would multiply (1 + tan x)*(1 + tan x) in the same way you would multiply (a+b)*(a+b). @dbbl21

OpenStudy (anonymous):

Thanks for your reply. I did that and simplified it to: tan^2 x + 2 tan x + 1, then simplified this to: (sin^2 x / cos^2 x) + 2(sin x / cos x) + 1 I cannot figure out where to go from here.

Directrix (directrix):

@dbbl21 There are several ways to write the answer. That is why we need the answer options posted so that we will have an idea of how to proceed.

OpenStudy (anonymous):

If I am understanding you correctly there are no answer options, just to simplify.

Directrix (directrix):

tan^2 x + 2 tan x + 1 would be where I'd stop, then.

OpenStudy (anonymous):

I managed to simplify it down to 2 cos x(sin^2 x) + 2 sin x(cos^2 x). Not sure if it is correct but it is all I can come up with for now. Thanks for your help.

Directrix (directrix):

How is this: 2 cos x(sin^2 x) + 2 sin x(cos^2 x) simpler than this: tan^2 x + 2 tan x + 1? I'll check and see if they are the same.

OpenStudy (anonymous):

I would use a trigonometric identity to do this tan^2x+1=sec^2x so it would be (sec^2x)^2 or sec^4x right?

Directrix (directrix):

@GABI_J I don't think that would work because the posted problem is: (1 + tan x)²  ≠ (1 + tan²x) in the same way that (a + b)² ≠ (a² + b²) (1 + tan²x)² does equal sec^4 x for the reasons you gave above. Lately, I have been wondering if the problem were supposed to be this: (1 + tan²x) in which case the simplification would be just sec²x x.

OpenStudy (anonymous):

yeah, my bad. As you said tan^2 x + 2 tan x + 1 but even on that I would preform one more step sec^2x+2tanx.

Directrix (directrix):

@GABI_J How do you get from here: tan^2 x + 2 tan x + 1 to here: sec^2x+2tanx ?

OpenStudy (anonymous):

b/c tan^2x=sec^2x-1 so tan^2x+1=sec^2x